Source power setting: Set Vgate=-1.65, yields AME_E =+0.60
Discovered copol sidelobes at 0.050m from boresight. Is only 5dB down from peak at tilt table = 45deg. Not seen at table = 0 deg. Tracked down to gap in the TKRAM coverage. Added another block of TKRAM to cover the gap. Sidelobe gone.
Failure to lock source. photomixer current needs to be -0.3 to be healthy, but it is only -0.11 mA. Fiber cable problem. Morgan fixed. Now getting -0.62mA. The receiver IFpower is 3.5.
Told Greg about a mechanical fix to hold the fiber cable metal shell securely to avoid cable flexure from moving the connector.
Scan parameters: 121x121 over span of 0.175x0.175m